OpenStudy (anonymous):

In a right triangle the hypotenuse is x+6 and the other sides are x and 8. Try to find x using the pythagorean theorem.

OpenStudy (anonymous):

do you know the theorem?

OpenStudy (anonymous):

a^2=b^2+c^2.

OpenStudy (anonymous):

The problem is the x+6.

OpenStudy (anonymous):

yes. where in this case a is the hypotenuse. (x+6)^2 = x^2 + 8^2 to expand (x+6)^2, you can use FOIL, which is an acronym which helps you multiply (x+6)(x+6). With habit, you'll see that (a+b)^2 is always a^2 + 2ab + b^2 so you can open it right away to x^2 + 12x + 36

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

glad i could help. lmk if u have follow-up questions ^_^